# Benchmarking Deep Learning-Based Methods for Irradiance Nowcasting with Sky Images

### v1

@@ 6 Conclusion
Due to the intermittent nature of solar PV energy, the growing integration of this energy source into the electrical grid presents several technical challenges, such as optimizing the operational cost of the grid and maintaining its frequency and voltage stability. To address this issue, recent studies have applied deep learning-based architectures and methods to forecast the short-term irradiance (or PV power) with the use of sky-image data, due to their high spatio-temporal resolution. Although most studies focus on the forecasting problem directly, this study prioritizes the nowcasting problem, which can be coupled with video prediction models to predict the future irradiance value. We benchmark 10 different model architectures and perform several ablation experiments to show that the choice of architecture can have large impacts on the quality of the predictions, as well as the choice of the data processing pipeline, such as applying a $\Delta\text{t}$ to the irradiance measurements, using a clear sky model to normalize the measurements, appending a Sun mask to the images, among others. We explored three different datasets that are commonly used by the solar forecasting community and show that our findings are consistent and can be validated across them all. In one of these datasets, we highlight a possible error related to the timestamps of the image data and provide a potential fix that we recommend readers use in future studies.

### v2

@@ 6 Conclusion
Due to the intermittent nature of solar PV energy, the growing integration of this energy source into the electrical grid presents several technical challenges, such as optimizing the operational cost of the grid and maintaining its frequency and voltage stability. To address this issue, recent studies have applied deep learning-based architectures and methods to forecast the short-term irradiance (i.e. nowcasting) from all-sky images. In this study, we approach the nowcasting problem using a two-step framework: first we predict the future ASI with a video prediction module; second we estimate the corresponding GHI to each predicted future image frame with an irradiance estimation module. Our primary focus is on the GHI estimation module, and we begin our analysis with a benchmark of 10 different model architectures. This benchmark revealed that the ResNet50 model was the best performer across all datasets and highlighted that model depth and residual connections can improve model performance, especially under cloudy sky conditions.
Using the ResNet50 model, we then conduct ablation experiments and uncover several key findings: 1) in the Folsom dataset, the image file names are misaligned with the irradiance timestamps, but the correct alignment can be achieved by using the image date modified metadata; 2) shifting the irradiance measurements by a small $\Delta\text{t}$ during training improves performance by mitigating the backward/forward averaging effects applied to these measurements; 3) predicting the clear-sky index $k_{t}$ yields substantial performance gains compared to predicting GHI directly; 4) appending a Sun mask as a fourth image channel only benefits the model performance when predicting the GHI directly; among other insights.
We conclude our study by combining our best GHI estimation model with a video prediction model, achieving state of the art performance on the SIRTA dataset and providing strong proof of concept results for this two-step approach. These results highlight the potential for significant advancements in both the video prediction and GHI estimation modules, and we hope they inspire further research using the two-step approach in the GHI nowcasting literature.
@@ Acknowledgements
The work of L. Varaschin was supported in part by Fundacao de Amparo à Pesquisa e Inovacao do Estado de Santa Catarina (FAPESC), Edital 61/2024.
This work was supported in part by computational resources funded by FAPESC under Grant 2024TR000090.